### v1

# AntibodyDesignBFN: High-Fidelity Fixed-Backbone Antibody Design via Discrete Bayesian Flow Networks

Abstract
The computational design of antibodies with high specificity and affinity is a cornerstone of modern therapeutic development. While deep generative models, particularly Denoising Diffusion Probabilistic Models (DDPMs), have demonstrated the ability to generate realistic antibody structures, they often suffer from high computational costs and the difficulty of modeling discrete variables like amino acid sequences. In this work, we present AntibodyDesignBFN, a novel framework for fixed-backbone antibody design based on Discrete Bayesian Flow Networks (BFN). Unlike standard diffusion models that rely on Gaussian noise removal or complex discrete corruption processes, BFNs operate directly on the parameters of the data distribution, enabling a continuous-time, fully differentiable generative process on the probability simplex. While recent pioneering works like IgCraft and AbBFN have introduced BFNs to the domain of antibody sequence generation and inpainting, our work focuses specifically on the inverse folding task—designing sequences that fold into a fixed 3D backbone. By integrating a lightweight Geometric Transformer utilizing Invariant Point Attention (IPA) and a resource-efficient training strategy with gradient accumulation, our model achieves superior performance. Evaluations on a rigorous 2025 temporal test set reveal that AntibodyDesignBFN achieves a robust 49.9% Overall Amino Acid Recovery (AAR), demonstrating that BFNs, when conditioned on 3D geometric constraints, offer a robust mathematical framework for high-fidelity antibody design. Code and model checkpoints are available at https://github.com/YueHuLab/AntibodyDesignBFN and https://huggingface.co/YueHuLab/AntibodyDesignBFN, respectively.
1 Introduction
Antibodies are critical components of the adaptive immune system and represent a dominant class of biotherapeutics. Their binding specificity is primarily governed by six complementarity-determining regions (CDRs) in conventional antibodies, or three CDRs on the heavy chain in the case of single-domain antibodies (nanobodies or VHHs). Among these, the third CDR of the heavy chain (H-CDR3) is typically the most diverse and functionally significant loop for antigen recognition. The ”inverse folding” problem in antibody engineering often entails designing specific amino acid sequences—particularly for the variable CDR loops—that adopt a desired 3D backbone conformation and bind a target epitope, while keeping the rest of the scaffold fixed.
Traditional approaches often rely on physics-based energy minimization (e.g., Rosetta [4]), which is computationally intensive and limited by the accuracy of energy functions. In recent years, deep learning methods have introduced generative paradigms, ranging from autoregressive models (e.g., ProteinMPNN [1]) to joint sequence-structure diffusion models like DiffAb [5] and AlphaPanda [3].
The emergence of Bayesian Flow Networks (BFNs) [2] has provided a powerful new alternative to diffusion models. BFNs reformulate generation as a Bayesian inference process, where the model iteratively updates its belief about the data distribution. IgCraft [7] and AbBFN [6] have recently demonstrated the power of BFNs for paired antibody sequence generation and motif scaffolding. However, these models primarily focus on sequence-space evolution or involve joint generation of multiple modalities including metadata, often without explicit structural conditioning on the antigen. Our work, AntibodyDesignBFN, extends this paradigm by conditioning the discrete BFN generative process on a full 3D atomic backbone. Crucially, this allows our model to explicitly ”see” the antigen epitope geometry during generation, ensuring that the designed sequences are not only naturally diverse but also structurally and electrostatically compatible with the target interface.
2 Methodology
2.1 Discrete Bayesian Flow Networks
We formulate the antibody sequence design problem using the Discrete Bayesian Flow Network framework. Let $\mathbf{x}$ be a ground-truth amino acid sequence of length $L$. In our fixed-backbone setting, the length $L$ is strictly determined by the input 3D structure. The Bayesian distribution is defined over the logits $\mathbf{\theta}\in\mathbb{R}^{L\times K}$ of the categorical distribution for the entire sequence simultaneously. This means the model updates its belief about every residue in parallel, capturing global dependencies without autoregressive bias.
The core distinction of BFNs is that they model the parameters (logits) of the sequence distribution rather than the sequence tokens themselves. This allows us to use continuous Gaussian processes to generate discrete data.
2.1.1 The Bayesian Sender (Forward Process)
The ”Sender” distribution $P_{S}(\mathbf{\theta}|\mathbf{x};t)$ defines how our information (belief) about the sequence evolves over continuous time $t\in[0,1]$. This is fundamentally different from Diffusion Models:
  • Diffusion Models: Corrupt the data $\mathbf{x}$ directly (e.g., by flipping tokens or adding Gaussian noise to embeddings).
  • BFNs: Transmit the true data $\mathbf{x}$ through a noisy channel into the parameter space $\mathbf{\theta}$.
The parameters $\mathbf{\theta}(t)$ are sampled from a Gaussian distribution:
$\mathbf{\theta}(t)\sim\mathcal{N}(\beta t\cdot\mathbf{e}_{\mathbf{x}},\beta t\cdot\mathbf{I})$ (1)
Here, $\beta$ is a signal-to-noise ratio parameter. At $t=0$ (Prior), the distribution $\text{Softmax}(\mathbf{\theta})$ is Uniform (Maximum Uncertainty). At $t=1$ (Data), the mean is $\beta\mathbf{e}_{\mathbf{x}}$, and the distribution is sharp (Maximum Certainty).
2.1.2 The Geometric Receiver (Architecture)
The ”Receiver” is a neural network $\Psi$ that estimates the true data distribution given the noisy parameters $\mathbf{\theta}$ and time $t$. The input to the network is the current ”belief” about the sequence, represented as $\mathbf{p}_{in}=\text{Softmax}(\mathbf{\theta})$. This soft probability distribution is concatenated with sinusoidal time embeddings and passed to the geometric encoder:
$\mathbf{p}_{out}=\Psi(\mathbf{p}_{in},t,\mathcal{G}_{backbone})$ (2)
We employ a Geometric Transformer utilizing Invariant Point Attention (IPA) layers. This explicitly conditions the sequence belief update on the 3D coordinates of the N, $C_{\alpha}$, C atoms. The attention mechanism calculates weights based on both the semantic similarity of the intermediate features and the Euclidean distance between residues in the rigid backbone frame. This ensures that residues spatially close in the 3D structure can communicate efficiently, even if they are distant in the primary sequence.
2.1.3 Continuous-Time Loss Function
The training objective simplifies elegantly to a weighted reconstruction loss:
$\mathcal{L}(\mathbf{x})=\beta\int_{0}^{1}\mathbb{E}_{\mathbf{\theta}\sim P_{S}(\cdot|\mathbf{x};t)}{||\mathbf{e}_{\mathbf{x}}-\hat{\mathbf{p}}_{out}(\mathbf{x}|\mathbf{\theta},t)||^{2}}dt$ (3)
In practice, we minimize a cross-entropy loss $\log(\mathbf{p}_{out}(\mathbf{x}_{true}))$, which is numerically more stable and optimizes the same objective.
2.2 Bayesian Sampling (Inference)
The generation process starts with a uniform prior $\mathbf{\theta}_{0}=\mathbf{0}$. At each step $k$ (of $N$ total steps), we perform a Bayesian update:
$\mathbf{\theta}_{k+1}=\underbrace{\mathbf{\theta}_{k}}_{\text{Prior}}+\underbrace{\frac{\beta}{N}\hat{\mathbf{p}}_{out}(\mathbf{\theta}_{k},t_{k})}_{\text{Evidence}}+\underbrace{\sqrt{\frac{\beta}{N}}\mathbf{\epsilon}_{k}}_{\text{Uncertainty}}$ (4)
This process can be viewed as an iterative refinement where the sequence emerges from the noise, guided by the structural constraints.
2.3 Training Strategy
We utilized gradient accumulation (accum_steps=8) to overcome memory limitations on consumer hardware (Mac mini M4), simulating a larger effective batch size. We employed FP16 mixed-precision training (via torch.cuda.amp or torch.mps) and a linear warmup strategy for the first 1,000 steps followed by a constant learning rate of 1e41e-4.
3 Experiments
3.1 Experimental Setup
Dataset: We utilized the Structural Antibody Database (SAbDab), filtered for non-redundancy (95% sequence identity clusters) with a training cutoff date of 2022.
Test Set: To rigorously evaluate the generalization capability of AntibodyDesignBFN and prevent any potential sidechain information leakage (a common pitfall in fixed-backbone design), we curated a novel test set consisting of 41 antibody-antigen complexes released in 2025, which were strictly excluded from our training data. Crucially, we pre-processed these structures by mutating all residues within the CDR loops to Poly-Alanine (PolyA), removing all native sidechain information. This forces the model to recover the sequence solely from the backbone geometry and the antigen context. All structures contain both the antibody Fv fragment and a bound antigen, ensuring that the fixed-backbone condition reflects real binding geometries. The PDB IDs are: 8S6T, 8S6V, 8UHP, 8UIG, 8UIH, 8UJI, 8V13, 8YYZ, 8Z2V, 8Z39, 8Z3A, 9C7X, 9CDS, 9CFD, 9CQA, 9CR9, 9D41, 9DBO, 9DQ3, 9DQ4, 9DST, 9EHT, 9EZE, 9FIK, 9GGP, 9GOX, 9H4R, 9J8A, 9JCY, 9JD1, 9M5B, 9MIC, 9MID, 9MIF, 9NJY, 9NPI, 9OAR, 9P4C, 9PI9, 9PWN, 9UK5.
3.2 Results
We evaluate performance using Amino Acid Recovery (AAR). To provide a rigorous baseline, we also evaluated ProteinMPNN [1] on the same 2025 dataset. Table 1 presents the comparative results using the finetuned model.
Table 1: Comparative Performance on the 2025 Temporal Test Set (41 Antibody-Antigen Complexes). The reported AAR values reflect the robustness of sequence recovery under fixed-backbone conditions.
Region ProteinMPNN AAR (%) Our AAR (%)
H-CDR1 50.7 68.6
H-CDR2 48.4 48.9
H-CDR3 36.8 27.4
L-CDR1 40.5 60.9
L-CDR2 46.0 50.2
L-CDR3 46.8 48.4
Average 44.2 49.9
4 Discussion
AntibodyDesignBFN distinguishes itself from existing protein design methodologies through its unique combination of Discrete Bayesian Flow Networks and explicit geometric conditioning. Compared to the widely used ProteinMPNN [1] baseline, our finetuned model achieves a significantly higher overall AAR of 49.9% (vs. 44.2%) on the 2025 temporal split—an improvement of 5.7 percentage points.
The performance gap is broadly consistent, with AntibodyDesignBFN outperforming ProteinMPNN on most CDR regions. Notably, our model achieves dramatic gains on H-CDR1 (68.6% vs. 50.7%) and L-CDR1 (60.9% vs. 40.5%), suggesting superior handling of canonical loop geometries. While ProteinMPNN shows higher recovery on H-CDR3 (36.8% vs. 27.4%), this region is known for its high variability and flexibility. Our model’s divergent designs in this region may reflect a stronger sensitivity to the antigen context (as evidenced by our antigen-impact analysis), leading it to design novel binders rather than strictly copying the native sequence. This domain adaptation capability suggests our model generates sequences that are biologically plausible and structurally compatible.
AI Usage Declaration
This manuscript was prepared with the assistance of large language models, specifically Gemini 2.0 and Google Antigravity, which were used for drafting text, refining mathematical formulations, checking LaTeX syntax, and exploring related literature. All content, data, and conclusions have been manually reviewed, verified, and authenticated by the authors, who take full responsibility for the accuracy, integrity, and originality of the work.